Problem 6: Frequencies of the string For a string stretched between two supports, two successive standing-wave frequencies are 510 Hz and 680 Hz. There are also other standing waves with higher and lower frequencies. What is the order of what is the length of the string?
Answer #2

SolcFion n v 2 Y85 2 L 510 - 2L 853408 5 Sb,
Answer #1

The expression for the frequency, "f(n)" of nth harmonic is given by –

f(n) = {(n+1)*V}/(2L) -------------------------------(i)

So –

f(n+1) = {(n+2)*V}/(2L) ----------------------------(ii)

Now do [(ii) – (i)] –

Means –

f(n+1) - f(n) = (V/2L)

We have given that –

f(n+1) = 680 Hz

f(n) = 510 Hz

So –

680 – 510 = (V/2L)

=> V / 2L = 170 Hz

Also given that, V = 85 m/s

So, 85 / 2L = 170 Hz

=> 2L = 85 / 170 = 0.50 m

=> L = 0.50/2 = 0.25 m

Hence, length of the string = 0.25 m

Now, put n = 0,1,2,3…….

Frequency of 1st harmonic = V/2L = 170 Hz

Frequency of 2nd harmonic = 2*V/2L = 2*170 = 340 Hz

Frequency of 3rd harmonic = 3*V/2L = 3*170 = 510 Hz

Frequency of 4th harmonic = 4*V/2L = 4*170 = 680 Hz

Therefore, the mentioned harmonics are 3rd and 4th harmonics.
Answer #3

aluれの,一 2L じ 2L An m-pl Igo S1o3 OY dles ous,从